For which values of a and b is the line b 2x + y = tangent to the parabola 2

y = ax

when x = 2? 


1
Expert's answer
2021-01-19T04:34:02-0500

"\text{The equation of the parabola :- } y=ax^2\\\;\\
\therefore \text{Slope of the parabola at point x=2 ,}\\
\left(\dfrac{dy}{dx}\right)_{x=2}=\left(2ax\right)_{x=2}=2a\times 2=4a"


"\text{Now equation of the line : }\\
 2x+y=b\\
\Rightarrow y=-2x+b\\
\;\\
\text{So the slope of the line }=-2\\\;\\
\text{Now equating slopes, we get,}\\
4a=-2\\
\text{or, } a=-\dfrac{1}{2}"


"\text{Now, at }x=2,\\
\text{Ordinate of parabola} = a\times2^2=\dfrac{-1}{2}\times2^2=-2\\
\text{and Ordinate of the line}=-2\times 2+b=b-4\\\;\\
\text{Now equating the ordinates, we get,}\\
b-4=-2\\
\text{or, }b=2\\\;\\
\therefore \text{For } a=-\dfrac{1}{2}\;\;\text{and}\;b=2, \;\;\;\; \text{2x+y=b is a tangent of $y=ax^2$.}"


So,


"\fcolorbox{black}{yellow}{\textcolor{blue}{$ a=-\dfrac{1}{2},\;b=2$}}"